# Predictive Analytics of Selections of Russet Potatoes

### v3

@@ Abstract
We explore the application of machine learning algorithms specifically to enhance the selection process of Russet potato clones in breeding trials by predicting their suitability for advancement. This study addresses the challenge of efficiently identifying high-yield, disease-resistant, and climate-resilient potato varieties that meet processing industry standards. Leveraging data from manually collected trials in the state of Oregon, we investigate the potential of a wide variety of state-of-the-art binary classification models. The dataset includes 1086 clones, with data on 38 attributes recorded for each clone, focusing on yield, size, appearance, and frying characteristics, with several control varieties planted consistently across four Oregon regions from 2013-2021. We conduct a comprehensive analysis of the dataset that includes preprocessing, feature engineering, and imputation to address missing values. We focus on several key metrics such as accuracy, F1-score, and Matthews correlation coefficient (MCC) for model evaluation. The top-performing models, namely a feedforward neural network classifier (Neural Net), histogram-based gradient boosting classifier (HGBC), and a support vector machine classifier (SVC), demonstrate consistent and significant results. To further validate our findings, we conducted a simulation study using the aims, data-generating mechanisms, estimands, methods, and performance measures (ADEMP) framework, simulating different data-generating scenarios to assess model robustness and performance through true positive, true negative, false positive, and false negative distributions, area under the receiver operating characteristic curve (ROC) and MCC. The simulation results highlight that non-linear models like SVC and HGBC consistently show higher ROC and MCC than logistic regression, thus outperforming the traditional linear model across various distributions, and emphasizing the importance of model selection and tuning in agricultural trials. Variable selection further enhances model performance and identifies influential features in predicting trial outcomes. The findings emphasize the potential of machine learning in streamlining the selection process for potato varieties, offering benefits such as increased efficiency, substantial cost savings, and judicious resource utilization. Our study contributes insights into precision agriculture and showcases the relevance of advanced technologies for informed decision-making in breeding programs.
@@ 1 Introduction
Potato (Solanum tuberosum L.) is an important food crop in the United States with an annual potato production of 392.24 million cwt in 2022. The Pacific Northwest forms the largest sub-region in potato production in the U.S., contributing 60.1% of U.S. potato production according to USDA’s National Agricultural Statistics Service. Potatoes are not only consumed as a fresh vegetable but also form a raw material for the French fry, chip, and starch processing industries. In the U.S., 39% of the potato crop is utilized by the frozen processing sector. The Columbia basin of Oregon and Washington states is the leader in the frozen potato sector, with more than 80% of potatoes grown here being russet type for utilization by the processing industry.
@@
The complexity further increases due to the intricate interplay between distinct traits and environmental interactions needed for multi year and multi location trials. Breeders must decide whether the potato clone should go into the subsequent selection trial stage based on its performance or whether retaining a poorly performing clone could add unnecessary cost to the program resulting in decreased breeding selection efficiency. Advances in genetic tools have resulted in the use of molecular markers and genomic selection to improve selection efficiency. Even with molecular and genomic data, it is essential to carry multi year, multi location trials and the breeder must select which clones to advance. In this context, artificial intelligence (AI) or machine learning technologies can help evaluate potato clones and assist in making such decisions (Dwivedi et al., 2021). The ascendancy of AI technologies in revolutionizing data analytics is evident, particularly within breeding programs. Researchers have used machine learning models to obtain highly accurate predictions of root yield based on the vegetative indices of the different growth stages of cassava (Manihot esculenta Crantz) (Selvaraj et al., 2020). With the potent integration of machine learning algorithms and chemical composition as predictors, researchers have successfully predicted consumer flavor perceptions of blueberries and tomatoes (Colantonio et al., 2022).
@@ 1.1 Related Work
Another study of potato employed the machine learning models of support vector machines (SVM) and multi-layer perceptrons (MLP) to classify potato varieties based on macroscopic observations and measurements for selection or de-selection during trials much like those used in our study (Ozguven et al., ). For each of these two classification techniques, four models were constructed—two for binary classification and two for 4-class classification. The results were reported in terms of the accuracy score (number of correct predictions) with the SVM method demonstrating higher accuracy than MLP in the 4-class models. It is unclear how many features Ozguven et al. considered or how they addressed missing data. Further, no parameter tuning was performed.
@@ 1.2 Contributions of This Study
    Comprehensive Model Comparison: We systematically compare a diverse set of classification models, including traditional approaches such as Naïve Bayes and decision trees, as well as more advanced techniques like multi-layer perceptron classifiers (MLPC), histogram-based gradient boosting classifiers (HGBC), and support vector machines (SVM). This thorough comparison highlights the strengths and weaknesses of each model within the context of an imbalanced agricultural dataset. By understanding these strengths and weaknesses, we aim to improve the accuracy of selecting potato clones with desirable traits such as yield, disease resistance, and processing quality, which are important for successful breeding outcomes.
@@ 2.2 Feature Engineering
Not all 38 aforementioned features were used in our analysis, and some additional features were included. Features with over 400 not applicable (NA) ratings were dropped: “flower color”, “vine size”, and “maturity”. The feature “bruise data” was also dropped as it contained mostly text description and was missing for most observations. The following new features were added:
@@ 2.3 Data Imputation
Once numerical variables were standardized, we hot encoded the categorical variables—a process where all categorical variables’ values are accounted for all observations, but for an observation’s given category, its value is encoded to 1 while the other non-applicable values are encoded to zero. Currently our framework does not handle imputation on categorical variables, so observation for which categorical values were missing would be dropped. However the only categorical variables in our dataset are “Trial Region” and “Year in Trial”, and there are no missing values for them.
@@ 2.5 Classification Models and Evaluation Metrics
    Quadratic discriminant analysis: This supervised learning technique computes discriminant functions for each class, which are then used to determine the decision boundary between each class. Assuming each class density has a multivariate Gaussian distribution and by comparing the log-ratio of two classes posteriors, we can obtain the quadratic discriminant functions
@@
    Adaptive boosting: This is a commonly used ensemble learning approach in which a sequence of weak learners (models that are slightly better than guessing randomly), such as small decision trees, is repeatedly applied to modified versions of the data (i.e., by changing the weights applied to the training observations). The predictions from all the weak classifiers are then combined to produce a majority vote prediction (Hastie et al., 2009).
@@
    Multi-layer perceptron (MLP): This is a feedforward neural network with at least three layers—an input layer, one or more hidden layers, and an output layer—in which neurons are fully connected with a nonlinear activation function. The MLP classifier is trained by backpropagation to optimize a loss function. The Broyden – Fletcher – Goldfarb – Shanno (BFGS) method, the optimization algorithm used in this framework, is an iterative method for solving unconstrained nonlinear optimization problems (Hinton, 1990).
@@
    Support vector machine (SVM) with a radial basis function kernel: A linear support vector machine (SVM) identifies the plane that maximizes the distance between itself and the closest set of points (specifically the feature vectors from a set of observations) that belong to the two distinct categories. Figure 2(a) shows a linearly separable dataset with a linear SVM decision boundary.
@@
    An ensemble of classifiers along with a meta-classifier: Also called a stacking classifier, two or more classifiers are applied to the datasets and the resulting classification predictions are used to train a meta-classifier (Tang et al., 2014). The stacking classifier used in this study has gradient boosting and SVM as its level-1 classifiers, and logistic regression as its level-2 (or meta-) classifier.
@@ 2.7.1 Aims
The primary aim of this simulation study was to assess the performance of a subset of classification models used in the overall study, whose goal was to predict whether a potato variety will graduate to the next set of trials (according to binary response variable: true_keeps used in supervised training). We focused on the top three performing models, and added logistic regression (linear model) for baseline comparison. Focusing on these models helped us keep computation manageable and enabled us to use a comprehensive cross validated grid-search approach for each trial.
@@
    Determine the best-performing model in terms of True Positives (TP), False Positives (FP), False Negatives (FN), True Negatives (TN), Area Under the Receiver Operating Characteristic Curve (ROC), and Matthew’s Correlation Coefficient (MCC).
@@ 2.7.5 Performance Measures
  • Area Under the Receiver Operating Characteristic Curve (ROC)
@@ 3 Results
Table 2 presents the MCCs for the classification models explored in this study on the imputed and non-imputed datasets. Columns containing more than 400 NA’s were dropped before imputation and data cleaning, so that both the imputed and non-imputed datasets contained the same set of features. The top performing classifiers (highlighted in gray) were the feedforward neural network classifier (Neural Net), histogram-based gradient boosting (HGBC), and support vector machine (SVC), as each obtained an MCC score above or very close to 0.5 in the cross-validation and test for both imputed and non-imputed datasets with high significance. We note that AdaBoosting achieved similar performance to the declared top performing models, however we will not discuss it further as we will focus on HGBC, which uses the same underlying principle of an ensemble of weak learners but it is a more modern algorithm allowing additional hyperparameters to be set (e.g., regularization).
@@
Table 2: Best test Mathews correlation coefficients (MCC) and their respective 95% confidence intervals (CI) for all the classification models considered in this study. Top performing models (highlighted in gray) were the feedforward neural network classifier (Neural Net), histogram-based gradient boosting (HGBC) and support vector machine (SVC), on datasets with both multiple imputation and without imputation. The results presented here are from models using all available features which have less than 400 NA’s. QDA stands for Quadratic Discriminant Analysis.
@@ 3.1 Top Performing Models
From among the large family of classification models considered (Section 2.5), three models performed consistently well according to their MCC scores, with both imputed and and non-imputed data. They were the feedforward neural network classifier (Neural Net), the histogram-based gradient boosting classifier (HGBC), and the support vector machine classifier (SVC). The results from these models without variable selection are highlighted in Table 2 in light gray, and their confusion matrices are presented in Figure 3.
@@
Not surprisingly, yield components emerged as the most important factors, with variables like ”Total Yield,” ”%RB,” ”Y1s,” ”%1s,” and ”Over 20oz” consistently selected across all models and datasets. Quality traits also played a substantial part in the classification process with features such as ”Specific Gravity,” ”Fry Color,” ”Hollow Heart,” and ”Brown Center” frequently chosen by the models. Environmental factors, particularly Growing Degree Days (GDD) variables, were also often included in the selected features.
Trial information proved to be another critical aspect, with ”Yr in trial” and ”Trial Region” consistently appearing in the selected variables. This indicates that the stage of the trial and the location where it’s conducted significantly influence the classification results. Interestingly, variables related to defects, such as ”Blackspot,” ”IBS,” and ”Vascular,” were more frequently selected in HGBC models, suggesting this particular algorithm may be more attuned to these quality issues.
@@ 3.2 Simulation Study Results
Table 5: Performance metrics for original and simulated datasets.
@@
The simulation study conducted here offered valuable insights into the performance and robustness of various machine learning algorithms across different data distributions. By comparing the original data performance with results from four simulated distributions, we were able to draw significant conclusions about each model’s strengths and weaknesses. The simulation study results (see Table 5) indicate that non-linear models, such as Support Vector Machine (SVM), Histogram-based Gradient Boosting Classifier (HGBC), and Neural Networks, generally outperform traditional linear models like Logistic Regression in terms of classification metrics such as True Positives (TP), False Positives (FP), True Negatives (TN), False Negatives (FN), area under the receiver operating characteristic curve (ROC score), and Matthews Correlation Coefficient (MCC). These findings highlight the importance of selecting appropriate machine learning models for agricultural trials, where data can be complex and non-linear relationships are prevalent.
Looking at the original data performance, we find that each model has its unique characteristics. The Neural Net achieved high true positives (136) but also had relatively high false positives (10), resulting in the second-highest ROC score (0.73242). The HGBC demonstrated excellent precision with the highest true positives (143) and lowest false positives (3), although it had the lowest ROC score (0.656393) on the original data. The SVM showed a balanced performance with high true positives (142), low false positives (4), and the highest MCC (Matthews Correlation Coefficient) of 0.506071, suggesting good overall performance. Logistic Regression achieved the highest ROC score (0.752055) but at the cost of the highest false positives (14), indicating a potential trade-off between sensitivity and specificity.
When we examine the performance across the simulated distributions (normal_uniform, normal_beta, gamma_uniform, and gamma_beta), several interesting patterns emerge. All models maintained relatively consistent performance across these distributions, which suggests a degree of robustness to different data generation processes. However, we observe that ROC scores for the simulated datasets generally decreased compared to those of the original data, with the Neural Net showing the most significant drop (from 0.73242 to around 0.59). This decrease hints at the possibility that the models may be slightly overfitted to the original data distribution.
@@
The HGBC, on the other hand, demonstrated consistent performance across simulations. It maintained high true positives (139–140) and low false positives (6–7), suggesting good precision and specificity. The stability of its ROC scores across different distributions (around 0.65–0.66) indicates robustness, making it a reliable choice when the underlying data distribution is uncertain.
@@
Logistic Regression showed relatively stable performance across simulations and maintained the highest ROC scores in simulated data (around 0.67–0.68), suggesting good discrimination ability. However, it consistently had higher false positives (12–13) compared to other models, indicating a potential bias towards positive predictions. This characteristic could be beneficial or detrimental depending on the specific requirements of the application.

### v4

@@ Abstract
We explore the application of machine learning algorithms specifically to enhance the selection process of Russet potato (Solanum tuberosum L.) clones in breeding trials by predicting their suitability for advancement. This study addresses the challenge of efficiently identifying high-yield, disease-resistant, and climate-resilient potato varieties that meet processing industry standards. Leveraging manually collected data from trials in the state of Oregon, we investigate the potential of a wide variety of state-of-the-art binary classification models. The dataset includes 1086 clones, with data on 38 attributes recorded for each clone, focusing on yield, size, appearance, and frying characteristics, with several control varieties planted consistently across four Oregon regions from 2013-2021. We conduct a comprehensive analysis of the dataset that includes preprocessing, feature engineering, and imputation to address missing values. We focus on several key metrics such as accuracy, F1-score, and Matthews correlation coefficient (MCC) for model evaluation. The top-performing models, namely a feedforward neural network classifier (Neural Net), histogram-based gradient boosting classifier (HGBC), and a support vector machine classifier (SVM), demonstrate consistent and significant results. To further validate our findings, we conducted a simulation study using the aims, data-generating mechanisms, estimands, methods, and performance measures (ADEMP) framework, simulating different data-generating scenarios to assess model robustness and performance through true positive, true negative, false positive, and false negative distributions, area under the receiver operating characteristic curve (AUC-ROC) and MCC. The simulation results highlight that non-linear models like SVM and HGBC consistently show higher AUC-ROC and MCC than logistic regression (LR), thus outperforming the traditional linear model across various distributions, and emphasizing the importance of model selection and tuning in agricultural trials. Variable selection further enhances model performance and identifies influential features in predicting trial outcomes. The findings emphasize the potential of machine learning in streamlining the selection process for potato varieties, offering benefits such as increased efficiency, substantial cost savings, and judicious resource utilization. Our study contributes insights into precision agriculture and showcases the relevance of advanced technologies for informed decision-making in breeding programs.
@@ 1 Introduction
Potato (Solanum tuberosum L.) is an important food crop in the United States with an annual potato production of $\approx$ 43.9 billion pounds in 2022. The Pacific Northwest forms the largest sub-region in potato production in the U.S., contributing 60.1% of U.S. potato production according to USDA’s National Agricultural Statistics Service. Potatoes are not only consumed as a fresh vegetable but also form a raw material for the French fry, chip, and starch processing industries. In the U.S., 39% of the potato crop is utilized by the frozen processing sector. The Columbia basin of Oregon and Washington states is the leader in the frozen potato sector, with more than 80% of potatoes grown here being russet type for utilization by the processing industry.
@@
The complexity further increases due to the intricate interplay between distinct traits and environmental interactions needed for multi year and multi location trials. Breeders must decide whether the potato clone should go into the subsequent selection trial stage based on its performance or whether retaining a poorly performing clone could add unnecessary cost to the program resulting in decreased breeding selection efficiency. Advances in genetic tools have resulted in the use of molecular markers and genomic selection to improve selection efficiency. Even with molecular and genomic data, it is essential to carry multi year, multi location trials and the breeder must select which clones to advance. In this context, machine learning technologies can help evaluate potato clones and assist in making such decisions (Dwivedi et al., 2021). The ascendancy of machine learning technologies in revolutionizing data analytics is evident, particularly within breeding programs. Researchers have used machine learning models to obtain highly accurate predictions of root yield based on the vegetative indices of the different growth stages of cassava (Manihot esculenta Crantz) (Selvaraj et al., 2020). With the potent integration of machine learning algorithms and chemical composition as predictors, researchers have successfully predicted consumer flavor perceptions of blueberries and tomatoes (Colantonio et al., 2022).
@@ 1.1 Related Work
Another study of potato employed the machine learning models of support vector machines (SVM) and multi-layer perceptrons (MLP, also known as Neural Nets) to classify potato varieties based on macroscopic observations and measurements for selection or de-selection during trials much like those used in our study (Ozguven et al., ). For each of these two classification techniques, four models were constructed—two for binary classification and two for 4-class classification. The results were reported in terms of the accuracy score (number of correct predictions) with the SVM method demonstrating higher accuracy than MLP in the 4-class models. It is unclear how many features Ozguven et al. considered or how they addressed missing data. Further, no parameter tuning was performed.
@@ 1.2 Contributions of This Study
    Comprehensive Model Comparison: We systematically compare a diverse set of classification models, including traditional approaches such as Naïve Bayes and decision trees, as well as more advanced techniques like multi-layer perceptron classifiers (referred to as Neural Nets), histogram-based gradient boosting classifiers (HGBC), and support vector machines (SVM). This thorough comparison highlights the strengths and weaknesses of each model within the context of an imbalanced agricultural dataset. By understanding these strengths and weaknesses, we aim to improve the accuracy of selecting potato clones with desirable traits such as yield, disease resistance, and processing quality, which are important for successful breeding outcomes.
@@ 2.2 Feature Engineering
Not all 38 aforementioned features were used in our analysis, and some additional features were included. Features with over 400 not available (NA) ratings were dropped: “flower color”, “vine size”, and “maturity”. The feature “bruise data” was also dropped as it contained mostly text description and was missing for most observations. The following new features were added:
@@ 2.3 Data Imputation
Once numerical variables were standardized, we hot encoded the categorical variables—a process where all categorical variables’ values are accounted for all observations, but for an observation’s given category, its value is encoded to 1 while the other non-available values are encoded to zero. Currently our framework does not handle imputation on categorical variables, so observation for which categorical values were missing would be dropped. However the only categorical variables in our dataset are “Trial Region” and “Year in Trial”, and there are no missing values for them.
@@ 2.5 Classification Models and Evaluation Metrics
    Quadratic discriminant analysis: QDA is a supervised learning technique that computes discriminant functions for each class, which are then used to determine the decision boundary between each class. Assuming each class density has a multivariate Gaussian distribution and by comparing the log-ratio of two classes posteriors, we can obtain the quadratic discriminant functions
@@
    Adaptive boosting: AdaBoosting is a commonly used ensemble learning approach in which a sequence of weak learners (models that are slightly better than guessing randomly), such as small decision trees, is repeatedly applied to modified versions of the data (i.e., by changing the weights applied to the training observations). The predictions from all the weak classifiers are then combined to produce a majority vote prediction (Hastie et al., 2009).
@@
    Multi-layer perceptron (Neural Net): This is a feedforward neural network with at least three layers—an input layer, one or more hidden layers, and an output layer—in which neurons are fully connected with a nonlinear activation function. The neural net classifier is trained by backpropagation to optimize a loss function. The Broyden – Fletcher – Goldfarb – Shanno (BFGS) method, the optimization algorithm used in this framework, is an iterative method for solving unconstrained nonlinear optimization problems (Hinton, 1990).
@@
    Support vector machine (SVM) with a radial basis function (RBF) kernel: A linear support vector machine (SVM) identifies the plane that maximizes the distance between itself and the closest set of points (specifically the feature vectors from a set of observations) that belong to the two distinct categories. Figure 2(a) shows a linearly separable dataset with a linear SVM decision boundary.
@@
    An ensemble of classifiers along with a meta-classifier: Also called a stacking classifier, two or more classifiers are applied to the datasets and the resulting classification predictions are used to train a meta-classifier (Tang et al., 2014). The stacking classifier used in this study has gradient boosting and SVM as its level-1 classifiers, and logistic regression (LR) as its level-2 (or meta-) classifier.
@@ 2.7.1 Aims
The primary aim of this simulation study was to assess the performance of a subset of classification models used in the overall study, whose goal was to predict whether a potato variety will graduate to the next set of trials (according to binary response variable: true_keeps used in supervised training). We focused on the top three performing models, and added logistic regression (linear model) for baseline comparison. Focusing on these models helped us keep computation manageable and enabled us to use a comprehensive cross validated grid search approach for each trial.
@@
    Determine the best-performing model in terms of True Positives (TP), False Positives (FP), False Negatives (FN), True Negatives (TN), Area Under the Receiver Operating Characteristic Curve (AUC-ROC), and Matthew’s Correlation Coefficient (MCC).
@@ 2.7.5 Performance Measures
  • Area Under the Receiver Operating Characteristic Curve (AUC-ROC)
@@ 3 Results
Table 2 presents the MCCs for the classification models explored in this study on the imputed and non-imputed datasets. Columns containing more than 400 NA’s were dropped before imputation and data cleaning, so that both the imputed and non-imputed datasets contained the same set of features. The top performing classifiers (highlighted in gray) were the feedforward neural network classifier (Neural Net), histogram-based gradient boosting (HGBC), and support vector machine (SVM), as each obtained an MCC score above or very close to 0.5 in the cross-validation and test for both imputed and non-imputed datasets with high significance. We note that AdaBoosting achieved similar performance to the declared top performing models, however we will not discuss it further as we will focus on HGBC, which uses the same underlying principle of an ensemble of weak learners but it is a more modern algorithm allowing additional hyperparameters to be set (e.g., regularization).
@@
Table 2: Best test Mathews correlation coefficients (MCC) and their respective 95% confidence intervals (CI) for all the classification models considered in this study. Top performing models (highlighted in gray) were the feedforward neural network classifier (Neural Net), histogram-based gradient boosting (HGBC) and support vector machine (SVM), on datasets with both multiple imputation and without imputation. The results presented here are from models using all available features which have less than 400 NA’s. QDA stands for Quadratic Discriminant Analysis.
@@ 3.1 Top Performing Models
From among the large family of classification models considered (Section 2.5), three models performed consistently well according to their MCC scores, with both imputed and and non-imputed data. They were the feedforward neural network classifier (Neural Net), the histogram-based gradient boosting classifier (HGBC), and the support vector machine classifier (SVM). The results from these models without variable selection are highlighted in Table 2 in light gray, and their confusion matrices are presented in Figure 3.
@@
Not surprisingly, yield components emerged as the most important factors, with variables like “Total Yield”, “RB”, “Yield No. 1s”, “1s”, and “Over 20 oz.” consistently selected across all models and datasets. Quality traits also played a substantial part in the classification process with features such as “Specific Gravity”, “Fry Color”, “Hollow Heart”, and “Brown Center” frequently chosen by the models. Environmental factors, particularly Growing Degree Days (GDD) variables, were also often included in the selected features.
Trial information proved to be another critical aspect, with “Year in trial” and “Trial Region” consistently appearing in the selected variables. This indicates that the stage of the trial and the location where it’s conducted significantly influence the classification results. Interestingly, variables related to defects, such as “Blackspot”, “Internal Brown Spot”, and “Vascular”, were more frequently selected in HGBC models, suggesting this particular algorithm may be more attuned to these quality issues.
@@ 3.2 Simulation Study Results
Table 5: Performance metrics for original and simulated datasets. Note that ROC refers to the AUC-ROC score.
@@
The simulation study conducted here offered valuable insights into the performance and robustness of various machine learning algorithms across different data distributions. By comparing the original data performance with results from four simulated distributions, we were able to draw significant conclusions about each model’s strengths and weaknesses. The simulation study results (see Table 5) indicate that non-linear models, such as Support Vector Machine (SVM), Histogram-based Gradient Boosting Classifier (HGBC), and Neural Networks, generally outperform traditional linear models like Logistic Regression in terms of classification metrics such as True Positives (TP), False Positives (FP), True Negatives (TN), False Negatives (FN), area under the receiver operating characteristic curve (AUC-ROC score), and Matthews Correlation Coefficient (MCC). These findings highlight the importance of selecting appropriate machine learning models for agricultural trials, where data can be complex and non-linear relationships are prevalent.
Looking at the original data performance, we find that each model has its unique characteristics. The Neural Net achieved high true positives (136) but also had relatively high false positives (10), resulting in the second-highest AUC-ROC score (0.73242). The HGBC demonstrated excellent precision with the highest true positives (143) and lowest false positives (3), although it had the lowest AUC-ROC score (0.656393) on the original data. The SVM showed a balanced performance with high true positives (142), low false positives (4), and the highest MCC (Matthews Correlation Coefficient) of 0.506071, suggesting good overall performance. Logistic Regression (LR) achieved the highest AUC-ROC score (0.752055) but at the cost of the highest false positives (14), indicating a potential trade-off between sensitivity and specificity.
When we examine the performance across the simulated distributions (normal_uniform, normal_beta, gamma_uniform, and gamma_beta), several interesting patterns emerge. All models maintained relatively consistent performance across these distributions, which suggests a degree of robustness to different data generation processes. However, we observe that the AUC-ROC scores for the simulated datasets generally decreased compared to those of the original data, with the Neural Net showing the most significant drop (from 0.73242 to around 0.59). This decrease hints at the possibility that the models may be slightly overfitted to the original data distribution.
@@
The HGBC, on the other hand, demonstrated consistent performance across simulations. It maintained high true positives (139–140) and low false positives (6–7), suggesting good precision and specificity. The stability of its AUC-ROC scores across different distributions (around 0.65–0.66) indicates robustness, making it a reliable choice when the underlying data distribution is uncertain.
@@
Logistic Regression (LR) showed relatively stable performance across simulations and maintained the highest AUC-ROC scores in simulated data (around 0.67–0.68), suggesting good discrimination ability. However, it consistently had higher false positives (12–13) compared to other models, indicating a potential bias towards positive predictions. This characteristic could be beneficial or detrimental depending on the specific requirements of the application.